Calculate the pH of a buffer solution that is 0.398 M in NHz (ammonia) and 0.325 M in NHBr. pH =
Answer #1

pH= 14- (pky + log file) (Salt) – IN HqBv] - 0.325 M [base] = [NH₂ ] = 0.398 m pky of NHg = 4,75 :. pH= 14- (1.75 +19 ()) =pH